# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 6. Color swap
From this section to the end of the paper, we will assume that the color map $c^{(i)}$ one each border of a two-color hive considered is
$c^{(i)}=(\underset{d\text{ times}}{\underbrace{1,\ldots,1}},\underset{n-d\text{ times}}{\underbrace{0,\ldots,0}},2,\underset{d\text{ times}}{\underbrace{1,\ldots,1}}).$
Such a two-color hive is called regular.
@@ 6.1. Gash propagation
Let $H$ be a dual hive. We define a gash to be a pair of edges $(e_{1},e_{2})\in E_{n}^{2}$ sharing one vertex such that $C(e_{1})=0$, $C(e_{2})=1$ and such that $e_{1}$ and $e_{2}$ are both of the same type with $h(e_{1})>h(e_{2})$. We call the type of a gash $g$ the common type of its two edges.
Note that this definition only depends on the color map $C$ of $H$. Let $g$ be a gash on $\partial T_{n}^{(2)}$ for some dual hive $H$. Then $g$ is necessarily of type $2$. East to $g$, there are only six possible configurations given in Figure 12. In this section, we show that such a gash $g$ can be moved across the color map $C$ using local moves until reaching configuration $(v)$ or $(vi)$ of Figure 12.
Figure 12. The six possible adjacent configurations to a gash of type $1$ shown in dashed edges. The same holds for a gash of other types up to rotations.
Definition 6.2 (Gash propagation in $(i)$, $(ii)$ and $(iii)$.).
Let $g=(e_{1},e_{2})$ be a gash and let $c$ be the configuration of the color map $C$ at $g$.
@@
    If $c$ is either $(i)$ or $(ii)$, the propagation of $g$ through $c$ is the gash $g^{\prime}=(e_{1}^{\prime},e_{2}^{\prime})$ given by the pair of edges having the same type as $(e_{1},e_{2})$.
@@
    If $c$ is $(iii)$, the two other pairs of edges of the same type in the triangle $(iii)$ can be taken as the propagation of $g$.
6.1.1. Propagation through configuration (iv)
The configuration (iv) does not give an immediate way to continue the gash propagation. In this section, we give a way to propagate any gash having such configuration adjacent to it.
Definition 6.3 (Openings and arrows).
Consider either two edges $(e_{1},e_{2})$ of the same color 00 or $1$ sharing a vertex and forming an angle of $\frac{2\pi}{3}$. We call this a 00 or $1$ opening depending on the color. There are only two possible completions of an opening as shown in Figure 13 for the case of a 00 opening. Note that in completions $(0.II)$ and $(1.II)$, the two new edges $(e_{3},e_{4})$ both of the same color as $(e_{1},e_{2})$ form another opening of the same color. From an opening, denote $r\geq 0$ the number of successive completions of type $(II)$ before the first completion of type $(I)$. The resulting configuration is called an arrow of color 00 or $1$ depending of any of its $r+1$ openings and $r$ is the length of this arrow.
Figure 13. First row from left to right : an 00 opening, completion $(0.I)$, completion $(0.II)$ and an 00 arrow of length $4$. The second row is the analog for color $1$.
Let us call a lozenge rigid if the inner edge is colored $m$. Remark that configuration $(iv)$ corresponds to two openings neighbors to a same rigid lozenge. Each of the arrows on these openings has some finite length, say $l_{1},l_{2}\geq 0$. The region delimited by the two arrows is completely determined and consists of $l_{1}\times l_{2}$ rigid lozenges, see Figure 14.
Figure 14. The region $R$ in between two arrows is fully determined and consists in rigid lozenges with the same orientation as in the arrows.
Once $R$ is filled with rigid lozenges, there are two new openings, one of each color at the end of the two arrows. We call a $(l_{1},l_{2})$ region the configuration where two openings are on consecutive sides of a region filled with rigid lozenges such that there are $l_{1}$, respectively $l_{2}$, edges of type $2$ colored 00, respectively of type 00 colored $1$. Note that the opposite side of configuration $(iv)$ from the gash can be seen as a $(1,1)$ region. When $(l_{1},l_{2})=(0,0)$, we define the region to be the configuration on the right in Figure 15.
Figure 15. A $(l_{1},l_{2})$ region (left) and a $(0,0)$ region (right).
Lemma 6.4 (Growing region procedure).
Consider a $(l_{1},l_{2})$ region for any $l_{1},l_{2}\geq 1$. Let $l_{1}^{\prime},l_{2}^{\prime}\geq 0$ be the length of the arrows from the openings of color 00 and $1$ respectively. Then, there is a unique color map completion which results in a $(l_{1}-1+l_{2}^{\prime},l_{2}-1+l_{1}^{\prime})$ region.
@@ 6.1.1. Propagation through configuration (iv)
Let us first assume that $(l_{1},l_{2},l^{\prime}_{1},l^{\prime}_{2})\neq(1,1,0,0)$. The reasoning is similar to the one before Figure 14. The arrow of color 00 creates an empty region $R_{0}$ similar to $R$ of size $(l_{1}-1)\times l^{\prime}_{1}$ which can only be filled using rigid lozenges because of its boundary conditions. Likewise, the arrow of color $1$ creates an empty region $R_{1}$ of size $(l_{2}-1)\times l^{\prime}_{2}$ which can also only be filled using rigid lozenges. The regions $R_{0}$ and $R_{1}$ define a region $R_{2}$ which by the same argument can only be filled using rigid lozenges. After filling every region $R_{i},0\leq i\leq 2$, we obtain a $(l_{1}-1+l_{2}^{\prime},l_{2}-1+l_{1}^{\prime})$ region. In the case where $(l_{1},l_{2},l^{\prime}_{1},l^{\prime}_{2})=(1,1,0,0)$, we obtain a $(0,0)$ region. ∎
Lemma 6.5 (General form of region $(iv)$).
Let $c$ be a configuration $(iv)$ where we view its east boundary as a $(1,1)$ region. There exists a finite sequence $\left((l_{1}^{(1)},l_{2}^{(1)}),(l_{1}^{(2)},l_{2}^{(2)}),\dots,(l_{1}^{(r)},l_{2}^{(r)})\right)$, $r\geq 1$ such that $(l_{1}^{(i+1)},l_{2}^{(i+1)})$ is obtained from $(l_{1}^{(i)},l_{2}^{(i)})$ by the growing procedure of Lemma 6.4 and such that $(l_{1}^{(r)},l_{2}^{(r)})=(0,0)$.
@@
Lemma 6.4 shows that at each step of the arrow growing procedure, the region obtained is also $(l_{1},l_{2})$ region. Since the hive is finite, the procedure has to terminate which only happens by a $(0,0)$ region. ∎
Definition 6.6 (Propagation through $(iv)$).
Let $g$ be a gash on a configuration $(iv)$ as in Figure 12. Let $R$ be the region obtained by the arrow growing procedure of Lemma 6.5. The propagation of $g$ is the gash $g^{\prime}$ which is constituted by the pair of edges of different colors sharing a vertex both of the same type as $g$ in the last $(0,0)$ region.
6.1.2. From a gash to a rigid lozenge
Since one know how to propagate a gash, we now define the following procedure.
Proposition 6.7 (Gash to rigid lozenge).
Let $H$ be a hive and assume that there exists a gash $g$ on $\partial_{\lambda}H$. Apply the following steps:
@@ 6.1.2. From a gash to a rigid lozenge
    Start from the south-most gash $g$ on $\partial T_{n}^{(2)}$.
@@
    Propagate $g$ via $(i),(ii)$ or $(iv)$ configurations. If the propagation encounters a configuration $(v)$ or $(vi)$, STOP. If the propagation encounters a configuration $t$ of type $(iii)$, apply $(2^{\prime})$:
  3. (2’)
    Propagate the gash of type 00 from south border of $t$ using $(i),(ii)$ or $(iv)$. If the propagation encounters configuration $(v)$ or $(vi)$, STOP. Otherwise, when arriving at downward $(iii)$ or $\partial T_{n}^{(0)}$, go to step $(3)$.
@@
    Consider the gash $g^{\prime}$ of type $2$ on the east border of $t$. Continue gash propagation from $g^{\prime}$ with the additional rule that whenever propagation meets a downward $(iii)$ the next gash in the propagation is the gash of type $1$ or $2$ on its east side.
The above algorithm terminates on a configuration that is either $(v)$ or $(vi)$ in $H$.
For the proof of Proposition 6.7, we will need the following lemma which shows that the propagation cannot end by a $(iii)$ and that it ends with a configuration after the last $(iii)$ met.
Lemma 6.8 (Last (iii) configuration).
Suppose a gash $g$ of type $1$ arrives on an upward $(iii)$ configuration $t$. Consider the resulting gash $g^{\prime}$ of type $2$ on the east border of $t$. If $g^{\prime}$ does not encounter any $(iii)$ configuration, then it ends at a $(v)$ or $(vi)$ configuration.
Before proving Lemma 6.8, we need Lemma 6.9 which shows that any triangular region having two of its sides with edges colored $1$ has all its edges colored $1$.
Lemma 6.9 (Regular equilateral triangles).
Let $C:E_{n}\rightarrow\{0,1,3,m\}$ be a color map on edges of $T_{n}$. Let $R$ be any subset of edges of $E_{n}$ such that $\partial R$ is an equilateral triangle of size $s\geq 1$. Assume that the south and west boundary edges of $R$ have color $C(e)=1$. Then, every edge in $R$ has color $1$.

### v2

@@ 6. Color swap
The goal of this combinatorial section is to exhibit a convex body of dimension $D=\frac{(n-1)(n-2)}{2}$ having integer points counted by quantum Littlewood-Richardson coefficients, so that the limit expression (3.26) converges to the volume of a polytope. From this section to the end of the paper, we set $k=n+d$ and $\xi=\mathrm{e}^{\frac{i\pi}{3}}$. We also assume that the color maps of dual hives are regular in the sense of Definition 6.1.
Definition 6.1 (Regular boundaries).
A color map $C:E_{k}\rightarrow\{0,1,3,m\}$ is regular (or has regular boundaries) if for every $i\in\{0,1,2\}$,
(6.1) $c^{(i)}=(\underset{d\text{ times}}{\underbrace{1,\ldots,1}},\underset{n-d\text{ times}}{\underbrace{0,\ldots,0}},\underset{d\text{ times}}{\underbrace{1,\ldots,1}}).$
Moreover, we say that a dual hive $H=(C,L)$ is regular if its color map $C$ is.
6.1. Arrows and hexagons
Let us start with some definitions on local configurations of edges in $E_{k}$.
Definition 6.2 (Openings and arrows).
Let $x\in T_{k}$. An opening of type $l\in\{0,1,2\}$ at $x$ is the pair of edges $(e,e^{\prime})\in E_{k}^{2}$ such that
$\{e,e^{\prime}\}$ =$=\{(x,x+\xi^{1+l}),(x,x+\xi^{5+l})\}\text{ or }\{e,e^{\prime}\}=\{(x+\xi^{2+l},x),(x+\xi^{4+l},x)\}$
$C(e)$ $=C(e^{\prime})\in\{0,1\}.$
The color of the opening is defined as the color of edges $e$ and $e^{\prime}$.
Consider an opening $a=(e,e^{\prime})$ at $x$ of type $l$ and color $c\in\{0,1\}$. Let $e^{\prime\prime}=e^{\prime\prime}(a)$ be the edge such that $e,e^{\prime}$ are edges of the lozenge with middle edge $e^{\prime\prime}$. The only possible colors of the edge $e^{\prime\prime}$ are $C(e^{\prime\prime})\in\{0,1\}$. If $C(e^{\prime\prime})=c$, faces of the lozenge with middle edge $e^{\prime\prime}$ have all of their edges colored $c$. If $C(e^{\prime\prime})\neq c$, then there is an opening $a^{\prime}$ of type $l$ and color $c$ at the other endpoint of $e^{\prime\prime}$. Note that there can only be finitely many such openings $r\geq 0$ before $C(e^{\prime\prime})=c$.
Definition 6.3 (Arrow).
Let $a=(e,e^{\prime})$ be an opening of type $l$ and color $c$. Let $r\geq 0$ be the number of successive openings having middle edge $e^{\prime\prime}$ such that $C(e^{\prime\prime})\neq c$ with $C(e^{\prime\prime})\in\{0,1\}$ as in the previous paragraph. We say that the configuration of edges consisting of the $r\geq 0$ successive pairs of $3$ and $m$ lozenges together with the pair of direct and reverse faces with boundary edges of color $c$ is an arrow of length $r\geq 0$ at the opening $a$.
See Figure 13 for examples of openings and arrows.
Figure 13. First row from left to right : an opening $a$ with color 00 and type 00 at $x$, the case $C(e^{\prime\prime})=c$, the case $C(e^{\prime\prime})\neq c$ and an arrow of length $r=4$. The second row is the analog for color $1$. Uncolored edges have color $3$.
Definition 6.4 (ABC hexagons).
Let $C$ be a color map and let $h$ be a hexagon, that is, the union of six triangular faces sharing one vertex in $T_{k}$. We say that $h$ is an ABC hexagon (for the color map $C$) if the color map $C$ restricted to $h$ is any of the three configurations in Figure 14 up to a rotation.
A rotation of an ABC hexagon $h$ is the replacement of the values of $C$ by the ones obtained from a rotation of $h$ which preserves the value of $C$ on the boundary $\partial h\cap E_{k}$.
Note that type B has three possible rotations whereas A and C only have two. For a hexagon $h$, denote $E_{h}$ the edges of $E_{k}$ which are in $h$.
Let $A$ be an arrow of length $r\geq 1$ and type $l$ at an opening with center $x$. The center of the last opening is $y\in\{x+r\xi^{l},x-r\xi^{l}\}$. Notice that if the color $c$ of $A$ is 00 (resp. $1$), then the hexagon $h(y)$ with center $y$ is of type $C$ (resp.) $A$. Applying a rotation to $h(y)$ yields an arrow $A^{\prime}$ of length $r-1$ of type $l$ at the same opening with center $x$. By applying hexagon rotations to $x+r\xi^{l},x+(r-1)\xi^{l},\dots,x+\xi^{l}$ (or $x-r\xi^{l},x-(r-1)\xi^{l},\dots,x-\xi^{l}$) in this order, one gets an arrow $R(A)$ of length $r$ of type $l$ the same opening with center $x+r\xi^{l}$. We call this sequence of $r$ hexagon rotations the reversal of the arrow $A$. An example of arrow reversal is given in Figure 15
Figure 15. Reversal of an arrow of length $4$ at $x$. Uncolored edges have color $3$.
@@ 6.2. Gash propagation
We define a local configuration called a gash in Definition 6.5 which one can propagate in a dual hive. Local propagation rules are given in Definition 6.6 and the general propagation algorithm on dual hives is defined in Definition 6.7. The goal of the propagation algorithm is to find rigid lozenges of a dual hive in view of the next section.
@@
Let $x\in T_{k}$. A gash $g$ with center $x=x(g)$ is the union of the two edges $(x,x-\xi^{2l}),(x+\xi^{2l},x)$ for $l\in\{0,1,2\}$ such that
[…]
The type of a gash denoted $t(g)$ is defined as the type $l\in\{0,1,2\}$ of its edges.
Note that this definition only depends on the color map $C$ of $H$. Let $g$ be a gash. There are only six possible configurations given in Figure 16 adjacent to $g$. In this section, we show that such a gash $g$ can be moved across the color map $C$ using local moves until reaching configuration $(v)$ or $(vi)$ of Figure 16.
Figure 16. The six possible adjacent configurations to a gash of type $2$ shown in dashed edges. The same holds for a gash of other types up to rotations.
Definition 6.6 (Gash propagation).
Let $g$ be gash of type $l$ with center $x$.
@@
    Suppose that $g$ is adjacent to a configuration $(i)$. Let $y=x+\xi^{4}$, (resp. $y=x+1$, $y=x+\xi^{5}$) if $l=0$, (resp. $l=1,l=2$). We call the propagation of $g$ the gash $g^{\prime}$ of type $l$ at center $y$.
@@
    Suppose that $g$ is adjacent to a configuration $(ii)$. Let $y=x+\xi^{5}$, (resp. $y=x+\xi,y=x+1$) if $l=0$, (resp. $l=1,l=2$). We call the propagation of $g$ the gash $g^{\prime}$ of type $l$ at center $y$.
  3. (3)
    Suppose that $g$ is adjacent to a configuration $(iii)$. Let $y=x+1$ if $l=1$ or $l=2$. We call the propagation of $g$ the gash $g^{\prime}$ of type $3-l$ at center $y$.
  4. (4)
    Suppose that $g$ is adjacent to a configuration $(iv)$. Notice that there is a 00 opening at $x$ and thus an arrow of color 00 at $x$ with type $l+1$. Reverting this arrow yields a configuration $(i)$ adjacent to $g$ and we define the propagation of $g$ to be the gash $g^{\prime}$ of type $l$ as in step $(1)$. See Figure 17 for an illustration.
Figure 17. Propagation of a gash $g$ adjacent to a configuration $(iv)$. The arrow of color 00 has been reversed yielding a configuration $(i)$ adjacent to $g$.
Remark that the only propagation in which the type of the gash changes is $(iii)$. We now give a general procedure using local propagations from Definition 6.6. This procedure starts from a gash and propagates it until reaching a configuration that is either $(v)$ or $(vi)$.
Definition 6.7 (Propagation algorithm).
Define the following algorithm.
  1. Input:
    A color map $C$ and a gash $g$ of type $l\in\{1,2\}$.
  2. (1)
    Set $g^{(0)}=g$, $x^{(0)}=x(g)$, $t^{(0)}=t(g)$.
  3. (2)
    WHILE $g^{(s)}$ is adjacent to $(i),(ii),(iii)$ or $(iv)$: set $g^{(s+1)}$ to be the propagation of $g^{(s)}$ with center $x^{(s+1)}$ and type $t^{(s+1)}$.
Proposition 6.8 (Propagation algorithm is correct).
Let $g$ be a gash of type $2$ in $T_{k}$. The propagation algorithm terminates at a gash $\tilde{g}$ adjacent to configuration of type $(v)$ or $(vi)$.
For the proof of Proposition 6.8, we need Lemma 6.9 which shows that any triangular region having two of its sides with edges colored $c\in\{0,1\}$ has all its edges colored $c$.
Lemma 6.9 (Regular equilateral triangles).
Let $C:E_{k}\rightarrow\{0,1,3,m\}$ be a color map on edges of $T_{k}$. Let $R$ be any subset of edges of $E_{k}$ such that $\partial R$ is an equilateral triangle of size $s\geq 1$. Let $c\in\{0,1\}$ and assume that two boundaries of $R$ have edges $e$ which are all colored $c$. Then, every edge in $R$ is colored $c$.
@@ 1. Input:
Let us first show a general fact about a shape described below that we call a cup. For $r\geq 1$, we call a cup of length $r$ and type $i$ the union of $r$ consecutive type $i$ edges together with one edge of type $i+1$, respectively of type $i-1$, forming an angle of $\frac{2\pi}{3}$ with type $i$ edge with maximal and minimal heights. See Figure 18 for an example. Suppose that edges of a cup are all colored $c\in\{0,1\}$. Let us show by induction on $r$ that the only possible color of edges in the convex hull of the cup is $c$.
Figure 18. A cup of length $r=8$ and type 00 with color $c=1$. Edges in the convex hull of the cup are dotted
Two adjacent edges of the cup of different types form an opening of color $c$. One checks that the arrow at this opening has length zero for otherwise all the $r$ edges of type $i$ would belong to some non-rigid lozenges with opposite edges of type $i$ (resp. $i-1$) colored $c$ (resp. $1-c$) which is incompatible with the color $c$ of the other opening of the cup. Thus, the arrow has zero length. Then, the rest of the convex hull is a cup of length $r-1$ which completes the proof of cup completion by induction.
Let us consider the region $R$ of the statement, and suppose that the boundary of type $i-1$ and $i+1$ of $R$ are colored $c$. The corner of the triangular face between boundaries of type $i-1$ and $i+1$ has all of its edges colored $c$, as two of them lie on the boundary of $R$. This induces a cup of length $1$ and type $i$ having edges of color $c$. By the previous reasoning, its only color completion consists of edges of color $c$. Each completion of a cup of size $r$ with $r<s$ yields a cup of size $r+1$ with edges colored $c$. Filling cups of sizes $11,2,\dots,s-$ 1 with edges of color $c$ fills then $R$ with edges colored $c$ and thus proves the claim. ∎
Corollary 6.10 (Corners of regular boundaries).
Let $C:E_{k}\rightarrow\{0,1,3,m\}$ be a regular color map. Then, the three equilateral regions of side length $d$ each containing an extremal vertex of $T_{k}$ have all of their edges colored $1$.
@@
Notice that the regularity of the color map $C$ implies that each such triangular region of side length $d$ has two boundaries which lie on $\partial T_{k}^{(0)}\cup\partial T_{k}^{(1)}\cup\partial T_{k}^{(2)}$ having edges colored $1$. Applying Lemma 6.9 yields the result. ∎
Proof of Proposition 6.8.
At each step of the gash propagation, one has that $x^{(s+1)}_{0}<x^{(s)}_{0}$ or $x^{(s+1)}_{1}>x^{(s)}_{1}$ and $t^{(s)}\in\{1,2\}$ which implies that the while loop terminates on a last gash $g^{(\infty)}$. Assume for the sake of contradiction that $g^{(\infty)}$ is not adjacent to a configuration $(v)$ or $(vi)$. Since $g^{(\infty)}$ is the last gash, it is of type $1$ with center $x^{(\infty)}\in\partial T_{k}^{(1)}$. As the boundary $\partial T_{k}^{(1)}$ is regular, the edge of color $1$ in $g^{(\infty)}$ is the edge with height $n$ so that $x^{(\infty)}_{1}=n$, see Figure 19. Moreover, by Corollary 6.10, the equilateral triangle $R_{d}\subset E_{k}$ of length $d$ having one of its boundaries between $x^{(\infty)}$ and $(n,0)$ has all of its edges colored $1$.
Consider the last $(iii)$ configuration encountered before reaching $\partial T_{k}^{(1)}$ having a gash $g^{\prime}$ of type $1$. Note that such a configuration exists as all other configurations preserve the type of the gash during propagation of Definition 6.6. Consider the last gash $g$ with center $x$ resulting from a propagation of type $(i)$ after $g^{\prime}$ with the convention that $g=g^{\prime}$ if no configuration $(i)$ or $(iv)$ happen after $g^{\prime}$ (recall that configuration $(iv)$ reduces to a propagation of type $(i)$, see step $(4)$ of Definition 6.6). Propagations $(g^{(s)},s\geq 0)$ after $g=g^{(0)}$ are then of type $(ii)$ so that for $s\geq 0$, $x^{(s+1)}_{1}=x^{(s)}_{1}$ and $x^{(s+1)}_{0}=x^{(s)}_{0}-1$. Since $x^{(\infty)}\in R_{d}$, we would have $x\in R_{d}$ and the edge $e^{\prime}\in E_{k}$ with origin $x+\xi^{5}$ of type 00 and color 00 in the last $(i)$ configuration before $g$ (or in the last configuration $(iii)$ if no such configuration $(i)$ exists) would have both its endpoints in $R_{d}$ which contradicts the fact that edges in $R_{d}$ are all colored $1$. See Figure 19 for an illustration of the above argument. ∎
Figure 19. The 00 colored edge $e^{\prime}$ in the last configuration $(i)$ would be in $R_{d}$.
Remark 6.11 (Type $1$ propagation algorithm).
Note that the argument given in the proof of Proposition 6.8 remains valid in the case of a type $1$ gash $g$ with center $x$ such that the type 00 edge with origin $x+\xi^{5}$ is colored 00.
6.3. Color swap path
In Proposition 6.8, we showed that any gash $g$ of type $2$ can be propagated to find a configuration $(v)$ or $(vi)$ having a rigid lozenge in it. In this section, we show that via hexagon rotations of Definition 6.4, one can bring this rigid lozenge at the location of $x=x(g)$.
Lemma 6.12 (Back propagation).
Let $g$ be gash with center $x$ adjacent to a configuration $(i,ii)$ or $(iii)$ and suppose that its propagation $g^{\prime}$ with center $x^{\prime}$ is adjacent to a configuration $(v)$ or $(vi)$. Then, the hexagon $h^{\prime}$ with center $x^{\prime}$ is an ABC hexagon. Moreover, using a hexagon rotation of $h^{\prime}$, the hexagon $h$ with center $x$ is an ABC hexagon.
Proof.
One can check that the statement of Lemma 6.12 holds in all possible cases of configurations, see Figure 14. ∎
Proposition 6.13 (Gash reduction).
Let $C$ be a color map and let $g$ be a gash of type $2$ in $C$. Let $\tilde{g}=g^{(s)}$ for some $s\geq 0$ be the last gash as in Proposition 6.8. Using hexagon rotations of ABC hexagons with centers given by $x^{(s)},x^{(s-1)},\dots,x^{(0)}$ in this order, $C$ can be mapped to a color map $C^{\prime}$ such that $g$ is a gash for $C^{\prime}$ adjacent to a configuration $(v)$ or $(vi)$.
Proof.
Applying Lemma 6.12 to every center $x^{(s)},x^{(s-1)},\dots,x^{(1)}$ in this order yields the desired configuration. ∎
Remark 6.14 (Reduction of a type $1$ gash).
As in Remark 6.11, the result of Proposition 6.13 still holds if one considers $g$ of type $1$ with center $x$ such that the type 00 edge with origin $x+\xi^{5}$ is colored 00.
6.4. Color map reduction
Recall that we only consider regular boundary conditions for $T_{k}$ that is, the color map $C:E_{k}\rightarrow\{0,1,3,m\}$ is given by $1\dots 10\dots 01\dots 1$ on every boundary of $T_{k}$, where there are $d$ ones on each side of the $n-d$ zeros. The goal of this section is to show that any regular color map can be mapped via hexagon rotations to the simple color map of Definition 6.15.
Definition 6.15 (Simple color map).
Let $n\geq d\geq 0$ and $k=n+d$. The color map $C_{0}:E_{k}\rightarrow\{0,1,3,m\}$ called the simple color map is defined by
@@ 6.4. Color map reduction
    $C_{0}$ is regular and thus $C_{0}(e)=1$ for every edge $e\in E_{k}$ in any corner equilateral triangle of side length $d$ in $T_{k}$ as in Corollary 6.10,
@@
    $C_{0}(e)=1$ for every edge $e\in E_{k}$ in the lozenge of side length $d$ in $T_{k}$ having outer vertices $n\xi,n\xi+d\xi,n\xi+d,n\xi+d\xi^{5}$,
@@
    $C_{0}(e)=0$ for every edge $e\in E_{k}$ in the equilateral triangle having outer vertices $d,n$ and $d+n\xi$,
  4. (4)
    $C_{0}(e)=m$, respectively $C_{0}(e)=3$ for every edge $e\in E_{k}$ of type 00 with origin $x$ such that $d\leq x_{0}\leq n-1$ and $1\leq x_{1}\leq d$, respectively for every edge $e\in E_{k}$ of type 00 with origin $y$ such that $10\leq y_{0}\leq d-$ 1 and $d+1\leq y_{1}\leq n$.
Figure 20 shows an example of the simple color map $C_{0}$ for $k=5$ and $d=2$.
Figure 20. The color map $C_{0}$ for $k=5$ and $d=2$. Uncolored edges have color $3$ (picture done with the module Knutson-Tao puzzles of Sage [35]).
Proposition 6.16 (Color map reduction).
Let $C:E_{k}\rightarrow\{0,1,3,m\}$ be a regular color map. Using hexagon rotations, one can map $C$ to $C_{0}$, where $C_{0}$ is the simple color map of Definition 6.15.